# Discrete dark matter with light Dirac neutrinos

### v1

@@ IV Dark matter
In this model, gauge singlet Dirac fermions $N_{2,3}$, gauge singlet scalars $\chi_{2,3}$ and $SU(2)$ doublet scalars $\eta_{2,3}$, remain odd under the residual $\mathcal{Z}_{2}$ symmetry such that the lightest among them becomes a viable dark matter candidate. When considering the fermionic dark matter scenario, $i.e.$, $M_{N}<M_{\eta}<M_{\chi}$, achieving the correct relic density becomes challenging because the mass $M_{N}$ is constrained to the multi-TeV range by the neutrino mass criteria. Fermion DM in scotogenic model was studied in several earlier works, see Ahriche et al. (2018); Borah et al. (2020a); Mahanta and Borah (2019) for example. As can be seen from these earlier works, fermion singlet DM relic relies primarily on Yukawa portal annihilations and coannihilations. Due to the involvement of the same Yukawa couplings in both type-I and scotogenic contributions to neutrino mass, we can not have Yukawa coupling in $\geq\mathcal{O}(1)$ regime as discussed in the previous section. After including coannihilation among fermion DM $N$ and $\mathcal{Z}_{2}$-odd scalars, correct relic abundance is found only for Case-A with $M_{N}\in[3000,5000]$ GeV, considering the parameter space in agreement with neutrino data. However, for Case-B ($M_{N}\in[6000,10000]$ GeV) and Case-C ($M_{N}\in[11400,14000]$ GeV), where $M_{N}$ is pushed to even larger values, it is not possible to achieve the correct relic density for the fermionic dark matter via thermal freeze-out. This is because the large $M_{N}$ value and insufficient (co)annihilation rates always lead to an overabundant relic density. Thermal relic of $N$ in Case-A is obtained due to smaller $M_{N}$ and large connihilation rate with $SU(2)$ doublet scalars via large Yukawa portal ($y_{i}\gg y^{\prime}_{i}$) interactions. Since doublet scalars have more efficient annihilation rates, it is easier to obtain correct relic of $N$ in Case-A compared to Case-B where $N$ has sizeable Yukawa with scalars which are predominantly singlets. While Case-C offers the possibility of non-thermal or freeze-in fermion singlet DM, we find it to be overproduced as well leaving only Case-A as viable for fermion DM scenario.
@@
In the left panel of Fig. 8, we show the variation of scalar DM relic with its mass. As shown in several earlier works on scalar DM, the (co)annihilations mediated by electroweak gauge bosons and the SM Higgs are responsible for the dip for intermediate mass, leaving two distinct mass ranges satisfying correct relic. The right panel of the same figure shows the spin-independent DM-nucleon scattering cross-section (mediated via Higgs portal) as a function of scalar DM mass for the points which satisfy relic abundance criteria. We used the package MicrOMEGAs Alguero et al. (2024) to calculate the relic density and DM-nucleon scattering cross-section after generating the model files using LanHEP Semenov (2016). The cyan shaded region is ruled out by the most stringent direct search constraints from LZ-2022 Aalbers et al. (2022) for heavier DM and CRESST-III Abdelhameed et al. (2019) and DarkSide-50 Agnes et al. (2018) for lighter DM. Clearly we can see that the viable DM mass range that satisfies correct relic density and direct detection rate ranges from around $500$ GeV to several TeV scale. Interestingly, future direct detection experiment Darwin Aalbers et al. (2016) can probe a part of this high mass region upto $M_{\rm DM}\sim 3$ TeV.
@@ V Observable $\Delta N_{\rm eff}$
where $\rho_{R},\rho_{\gamma}$ denote total radiation and photon densities respectively. Existing data from CMB experiments like PLANCK constrains such additional light species by putting limits on the effective degrees of freedom for neutrinos during the era of recombination ($z\sim 1100$) as Aghanim et al. (2020)
@@
at $2\sigma$ or $95\%$ C.L. including baryon acoustic oscillation (BAO) data. At $1\sigma$ C.L. it becomes more stringent to ${\rm N}_{\rm eff}=2.99\pm 0.17$. Similar bound also exists from big bang nucleosynthesis (BBN), $3.42.3<{\rm N}_{\rm eff}<$ 3.4 at $95\%$ C.L. Cyburt et al. (2016). Recent observations by the DESI collaboration report $\Delta N_{\rm eff}<0.4$ Adame et al. (2024), which is weaker than the PLANCK bound given in Eq. (14). All these bounds are consistent with SM predictions ${\rm N^{SM}_{eff}}=3.045$ Mangano et al. (2005); Grohs et al. (2016); de Salas and Pastor (2016). Future CMB experiment CMB-S4 is expected reach a much better sensitivity of $\Delta{\rm N}_{\rm eff}={\rm N}_{\rm eff}-{\rm N}^{\rm SM}_{\rm eff}=0.06$ Abazajian et al. (2019), taking it closer to the SM prediction. Another future CMB experiment, specially, CMB-HD Aiola et al. (2022) can probe $\Delta N_{\rm eff}$ upto $0.014$ at $1\sigma$. Dirac nature of light neutrinos introduces additional relativistic degrees of freedom in the form of right chiral neutrinos which can contribute to $\Delta{\rm N}_{\rm eff}$ either thermally or non-thermally. Light Dirac neutrino models often lead to enhanced $\Delta{\rm N}_{\rm eff}$, some recent works on which can be found in Abazajian and Heeck (2019); Fileviez Pérez et al. (2019); Nanda and Borah (2019); Han et al. (2020); Luo et al. (2020a); Borah et al. (2020b); Adshead et al. (2020); Luo et al. (2020b); Mahanta and Borah (2021); Du and Yu (2021); Biswas et al. (2021); Borah et al. (2022a); Li et al. (2023); Das and Borah (2024). Depending upon the strength of Yukawa couplings associated with $\nu_{R}$, we can have either thermal or non-thermal contribution to $\Delta{\rm N}_{\rm eff}$ as we discuss in the following subsections.
@@ V.2 Non-thermal $\Delta N_{\rm eff}$
With $\chi$ in equilibrium and subsequently producing $N$ and $\nu_{R}$, the relevant Boltzmann equation for the comoving density of $\nu_{R}$ of flavour $\alpha$ is given by Biswas et al. (2023),
$\frac{d\tilde{Y}_{\nu^{\alpha}_{R}}}{dx}=\frac{\beta}{x\mathcal{H}s^{1/3}}\langle E\Gamma\rangle Y^{\rm eq}_{N}.$ (15)
Here, $x=\frac{M_{\chi}}{T}$, $\mathcal{H}$ is the Hubble parameter in the radiation dominated era and $\beta$ is defined as $\beta=\big{[}1+T\frac{dg_{s}/dT}{3g_{s}}\big{]}$ with $g_{s}$ being the effective relativistic entropy degrees of freedom. For the sub-eV scale $\nu_{R}$, which remains relativistic till recombination, the comoving energy density is defined as $\tilde{Y}_{\nu^{\alpha}_{R}}=\rho_{\nu^{\alpha}_{R}}/s^{4/3}$ with $\rho_{\nu^{\alpha}_{R}},s$ denoting energy density of $\nu^{\alpha}_{R}$ and entropy density of the Universe respectively. The quantity $\langle E\Gamma\rangle$ is defined as
$\langle E\Gamma\rangle=g_{N}g_{\nu_{R}}\frac{\lvert\mathcal{M}\rvert^{2}_{\chi\to\bar{\nu}^{\alpha}_{R}N}}{32\pi}\frac{(M^{2}_{\chi}-M^{2}_{N})^{2}}{M^{4}_{\chi}}.$ (16)
with $g_{i}$ being the internal degrees of freedom of species "$i$" and $|\mathcal{M}|_{\chi\to\bar{\nu}^{\alpha}_{R}N}$ represents the decay amplitude. The $\nu_{R}$ produced from $\chi$ decay will contribute to this additional radiation density. Therefore, with $\rho_{R}=\rho_{\rm SM}+\rho_{\nu_{R}}$, $\Delta N_{\rm eff}$ for one species of $\nu_{R}$ (indicated by flavour index $\alpha$) calculated at the CMB epoch $(T_{\rm CMB})$ is
@@
where, the extrapolation from the BBN epoch $(T_{\rm BBN})$ to CMB epoch $(T_{\rm CMB})$ is justified by the fact that both $\rho_{\nu_{L}},\rho_{\nu_{R}}$ redshift similarly for $T<T_{\rm BBN}$. One can simply calculate the energy density ratio at BBN epoch $\sim\mathcal{O}({10})$ MeV as $\nu_{R}$ production stops much earlier. The numerical factor 2 in Eq. (17) indicate the degrees of freedom of each $\nu_{R}$. The net contribution to $\Delta N_{\rm eff}$ can be obtained by summing over flavour index $\alpha$ as $\Delta N_{\rm eff}=\sum_{\alpha}\Delta N^{\alpha}_{\rm eff}$.
@@
The decoupling profile for $\nu_{R}$ for two different benchmark choices of Yukawa coupling $y_{i}^{\prime}$ is shown in Fig. 9. For scalar ($\chi$) and fermion ($N$) masses in the TeV ballpark, we find that for $y_{i}^{\prime}\leq\mathcal{O}(10^{-3})$, the interaction $N_{L}\bar{\nu_{R}}\leftrightarrow N_{L}\bar{\nu_{R}}$ can not attain thermal equilibrium in the early Universe. Therefore, Case-A with $y_{i}^{\prime}\sim\mathcal{O}(10^{-11})$, $\nu_{R}$ will never enter the thermal bath, hence, thermal $\Delta N_{\rm eff}$ is not possible. On the other hand, $\nu_{R}$ produced non-thermally from the decay of either $\chi$ or $N$ gives rise to negligible contribution to $\Delta N_{\rm eff}$. As mentioned before, the Case-B, with $0.01\leq y_{1}^{\prime}\leq 0.1$, $0.03\leq y_{2}^{\prime}\leq 0.1$, $0.03\leq y_{3}^{\prime}\leq 0.1$, we have thermal contribution to $\Delta N_{\rm eff}$. Due to decoupling of $\nu_{R}$ much above the electroweak scale, we have a fixed contribution $\Delta N_{\rm eff}=0.14$ for Case-B. Finally, in Case-C, Yukawa couplings are of the size $\mathcal{O}(10^{-6})$, and the corresponding interactions do not allow $\nu_{R}$ to enter the thermal bath. Considering the decay of $\chi\rightarrow\nu_{R}N$, we have shown the non-thermal contributions to $\Delta N_{\rm eff}$ in Fig. 10. Clearly, some part of the parameter space is already ruled out by PLANCK 2018 and DESI 2024 limits while most of the remaining parameter space remains within reach of future CMB experiments like SPT-3G, CMB-S4 and CMB-HD. Similar results can be obtained by considering the decay mode $N\rightarrow\nu_{R}\chi$ applicable for scalar DM, without changing the conclusion significantly.
@@ VI Other detection prospects
Firstly, we can have a lepton flavour violating (LFV) process at one-loop level, as shown in Fig. 11. The corresponding decay width can be calculated using the prescriptions given in Kuno and Okada (2001); Guo and Han (2020). With relevant Yukawa couplings of $y_{i}\sim\mathcal{O}(10^{-1}-10^{-2})$, as in Case-A, it gives the branching ratio $\text{Br}(\mu\rightarrow e\gamma)\sim\mathcal{O}(10^{-15})$, which is close to the recent upper bound indicated by MEG-II experiment $\text{Br}(\mu\rightarrow e\gamma)<3.1\times 10^{-13}$ Afanaciev et al. (2024). For Case-B and Case-C with smaller Yukawa couplings $y_{i}\sim\mathcal{O}(10^{-11})$ and $y_{i}\sim\mathcal{O}(10^{-7})$, we get very small branching ratio $\text{Br}(\mu\rightarrow e\gamma)$ as $\mathcal{O}(10^{-57})$ and $\mathcal{O}(10^{-41})$ respectively, beyond experimental reach in near future.
@@
Due to the presence of several new scalars, admixtures of singlet and doublet, we can have different collider prospects as well. If scalar DM mass $(M_{\chi})$ is lighter than the SM Higgs mass $(M_{h})$ such that $h\rightarrow\chi\chi$ decay is allowed, it can contribute to invisible Higgs decay. Such invisible Higgs decay is currently constrained at Br$(h\rightarrow{\rm invisible})<0.14$ Aad et al. (2022). Production of $\mathcal{Z}_{2}$-odd scalars in colliders like the large hadron collider (LHC) can give rise to pure leptonic final states plus missing transverse energy (MET) Gustafsson et al. (2012); Datta et al. (2017), hadronic final states plus MET Poulose et al. (2017) or a mixture of both. Such MET can arise from light Dirac neutrinos or scalar/fermion DM in our scenario. It is also possible to have more exotic signatures like tri-lepton plus MET Hashemi and Najjari (2016) and mono-jet Belyaev et al. (2016, 2019). It should be noted that we also have new scalars in the model which are $\mathcal{Z}_{2}$-even. They can lead to additional final states, similar to generic two Higgs doublet models. The new charged scalars can also modify the diphoton decay width of the SM Higgs, currently constrained at $\frac{{\rm BR}(h\rightarrow\gamma\gamma)_{\rm expt}}{{\rm BR}(h\rightarrow\gamma\gamma)_{\rm SM}}=1.12\pm 0.09$ Sirunyan et al. (2021).
@@
Another interesting detection prospect arises in the gravitational wave (GW) frontier. Spontaneous breaking of discrete symmetries like $A_{4}\times Z_{2}$, as we have in our model, can lead to formation of topological defects like domain walls (DW). While DW, if stable or allowed to dominate the energy density of the Universe, will be in conflict with cosmological observations, they can be made to disappear by introducing small bias, in the form of explicit symmetry breaking terms. Such annihilating or collapsing DW then generates stochastic GW which can be probed at ongoing or future experiments. GW signatures of minimal Dirac seesaw model was studied in earlier works Barman et al. (2022, 2023) while similar studies in the context of flavour symmetric models like $A_{4}$ can be found in Gelmini et al. (2021). The bias or explicit symmetry breaking terms introduced for DW disappearance may also lead to interesting indirect detection signatures of DM due to its decay into different standard model particles [^1] See King et al. (2024); Borah et al. (2024b) for related discussions.. While the model has several detection prospects, it can also be falsified by future observation of neutrinoless double beta decay.

### v2

@@ IV Dark matter
In this model, gauge singlet Dirac fermions $N_{2,3}$, gauge singlet scalars $\chi_{2,3}$ and $SU(2)$ doublet scalars $\eta_{2,3}$, remain odd under the residual $\mathcal{Z}_{2}$ symmetry such that the lightest among them becomes a viable dark matter candidate. When considering the fermionic dark matter scenario, $i.e.$, $M_{N}<M_{\eta}<M_{\chi}$, achieving the correct relic density becomes challenging because the mass $M_{N}$ is constrained to the multi-TeV range by the neutrino mass criteria. Fermion DM in scotogenic model was studied in several earlier works, see Ahriche:2017iar ; Borah:2018smz ; Mahanta:2019gfe for example. As can be seen from these earlier works, fermion singlet DM relic relies primarily on Yukawa portal annihilations and coannihilations. Due to the involvement of the same Yukawa couplings in both type-I and scotogenic contributions to neutrino mass, we can not have Yukawa coupling in $\geq\mathcal{O}(1)$ regime as discussed in the previous section. After including coannihilation among fermion DM $N$ and $\mathcal{Z}_{2}$-odd scalars, correct relic abundance is found only for Case-A with $M_{N}\in[3000,5000]$ GeV, considering the parameter space in agreement with neutrino data. However, for Case-B ($M_{N}\in[6000,10000]$ GeV) and Case-C ($M_{N}\in[11400,14000]$ GeV), where $M_{N}$ is pushed to even larger values, it is not possible to achieve the correct relic density for the fermionic dark matter via thermal freeze-out. This is because the large $M_{N}$ value and insufficient (co)annihilation rates always lead to an overabundant relic density. Thermal relic of $N$ in Case-A is obtained due to smaller $M_{N}$ and large co-annihilation rate with $SU(2)$ doublet scalars via large Yukawa portal ($y_{i}\gg y^{\prime}_{i}$) interactions. Since doublet scalars have more efficient annihilation rates, it is easier to obtain correct relic of $N$ in Case-A compared to Case-B where $N$ has sizeable Yukawa with scalars which are predominantly singlets. While Case-C offers the possibility of non-thermal or freeze-in fermion singlet DM, we find it to be overproduced as well leaving only Case-A as viable for fermion DM scenario.
@@
In the left panel of Fig. 8, we show the variation of scalar DM relic with its mass. As shown in several earlier works on scalar DM, the (co)annihilations mediated by electroweak gauge bosons and the SM Higgs are responsible for the dip for intermediate mass, leaving two distinct mass ranges satisfying correct relic. The right panel of the same figure shows the spin-independent DM-nucleon scattering cross-section (mediated via Higgs portal) as a function of scalar DM mass for the points which satisfy relic abundance criteria. We used the package MicrOMEGAs Alguero:2023zol to calculate the relic density and DM-nucleon scattering cross-section after generating the model files using LanHEP Semenov:2014rea . The cyan shaded region is ruled out by the most stringent direct search constraints from LZ-2022 LUX-ZEPLIN:2022qhg for heavier DM and CRESST-III CRESST:2019jnq and DarkSide-50 DarkSide:2018bpj for lighter DM. Clearly we can see that the viable DM mass range that satisfies correct relic density and direct detection rate ranges from around $500$ GeV to several TeV scale. Interestingly, future direct detection experiment Darwin DARWIN:2016hyl can probe a part of this high mass region up to $M_{\rm DM}\sim 3$ TeV.
@@ V Observable $\Delta N_{\rm eff}$
where $\rho_{R},\rho_{\gamma}$ denote total radiation and photon densities respectively. Existing data from CMB experiments like PLANCK constrains such additional light species by putting limits on the effective degrees of freedom for neutrinos during the era of recombination ($z\sim 1100$) as Planck:2018vyg
@@
at $2\sigma$ or $95\%$ C.L. including baryon acoustic oscillation (BAO) data. At $1\sigma$ C.L. it becomes more stringent to ${\rm N}_{\rm eff}=2.99\pm 0.17$. Similar bound also exists from big bang nucleosynthesis (BBN), $3.42.3<{\rm N}_{\rm eff}<$ 3.4 at $95\%$ C.L. Cyburt:2015mya . Recent observations by the DESI collaboration report $\Delta N_{\rm eff}<0.4$ at $2\sigma$ DESI:2024mwx , which is weaker than the PLANCK bound given in Eq. (14). All these bounds are consistent with SM predictions ${\rm N^{SM}_{eff}}=3.045$ Mangano:2005cc ; Grohs:2015tfy ; deSalas:2016ztq . Future CMB experiment CMB-S4 is expected reach a much better sensitivity of $\Delta{\rm N}_{\rm eff}={\rm N}_{\rm eff}-{\rm N}^{\rm SM}_{\rm eff}=0.06$ Abazajian:2019eic , taking it closer to the SM prediction. Another future CMB experiment, specially, CMB-HD CMB-HD:2022bsz can probe $\Delta N_{\rm eff}$ up to $0.014$ at $1\sigma$. Dirac nature of light neutrinos introduces additional relativistic degrees of freedom in the form of right chiral neutrinos which can contribute to $\Delta{\rm N}_{\rm eff}$ either thermally or non-thermally. Light Dirac neutrino models often lead to enhanced $\Delta{\rm N}_{\rm eff}$, some recent works on which can be found in Abazajian:2019oqj ; FileviezPerez:2019cyn ; Nanda:2019nqy ; Han:2020oet ; Luo:2020sho ; Borah:2020boy ; Adshead:2020ekg ; Luo:2020fdt ; Mahanta:2021plx ; Du:2021idh ; Biswas:2021kio ; Borah:2022obi ; Li:2022yna ; Das:2023yhv . Depending upon the strength of Yukawa couplings associated with $\nu_{R}$, we can have either thermal or non-thermal contribution to $\Delta{\rm N}_{\rm eff}$ as we discuss in the following subsections.
@@ V.2 Non-thermal $\Delta N_{\rm eff}$
With $\chi$ in equilibrium and subsequently producing $N$ and $\nu_{R}$, the relevant Boltzmann equation for the comoving density of $\nu_{R}$ of flavour $\alpha$ is given by Biswas:2022vkq ,
$\frac{d\tilde{Y}_{\nu^{\alpha}_{R}}}{dx}=\frac{\beta}{x\mathcal{H}s^{1/3}}\langle E\Gamma\rangle Y^{\rm eq}_{\chi}.$ (15)
Here, $x=\frac{M_{\chi}}{T}$, $\mathcal{H}$ is the Hubble parameter in the radiation dominated era and $\beta$ is defined as $\beta=\big{[}1+T\frac{dg_{s}/dT}{3g_{s}}\big{]}$ with $g_{s}$ being the effective relativistic entropy degrees of freedom. For the sub-eV scale $\nu_{R}$, which remains relativistic till recombination, the comoving energy density is defined as $\tilde{Y}_{\nu^{\alpha}_{R}}=\rho_{\nu^{\alpha}_{R}}/s^{4/3}$ with $\rho_{\nu^{\alpha}_{R}},s$ denoting energy density of $\nu^{\alpha}_{R}$ and entropy density of the Universe respectively. On the right hand side of Eq. (15), $Y^{\rm eq}_{\chi}\equiv n^{\rm eq}_{\chi}/s$ denotes the equilibrium comoving density of $\chi$. The quantity $\langle E\Gamma\rangle$ is defined as
$\langle E\Gamma\rangle=\frac{M_{\chi}}{2}\left(1-\frac{M^{2}_{N}}{M^{2}_{\chi}}\right)\Gamma_{\chi}$ (16)
where $\Gamma_{\chi}$ denotes the partial decay width for the process $\chi\rightarrow N\nu_{R}$. The $\nu_{R}$ produced from $\chi$ decay will contribute to this additional radiation density. Therefore, with $\rho_{R}=\rho_{\rm SM}+\rho_{\nu_{R}}$, $\Delta N_{\rm eff}$ for one species of $\nu_{R}$ (indicated by flavour index $\alpha$) calculated at the CMB epoch $(T_{\rm CMB})$ is
@@
where, the extrapolation from the BBN epoch $(T_{\rm BBN})$ to CMB epoch $(T_{\rm CMB})$ is justified by the fact that both $\rho_{\nu_{L}},\rho_{\nu_{R}}$ redshift similarly for $T<T_{\rm BBN}$. One can simply calculate the energy density ratio at BBN epoch $\sim\mathcal{O}({10})$ MeV as $\nu_{R}$ production stops much earlier. The numerical factor 2 in Eq. (17) indicate the degrees of freedom of each $\nu_{R}$. The net contribution to $\Delta N_{\rm eff}$ can be obtained by summing over flavour index $\alpha$ as $\Delta N_{\rm eff}=\sum_{\alpha}\Delta N^{\alpha}_{\rm eff}$. Ignoring the temperature dependence of the relativistic degrees of freedom, one can solve Eq. (15) analytically to find $\Delta N_{\rm eff}$ Biswas:2022vkq .
@@
The decoupling profile for $\nu_{R}$ for two different benchmark choices of Yukawa coupling $y_{i}^{\prime}$ is shown in Fig. 9. For scalar ($\chi$) and fermion ($N$) masses in the TeV ballpark, we find that for $y_{i}^{\prime}\leq\mathcal{O}(10^{-3})$, the interaction $N_{L}\bar{\nu_{R}}\leftrightarrow N_{L}\bar{\nu_{R}}$ can not attain thermal equilibrium in the early Universe. Therefore, Case-A with $y_{i}^{\prime}\sim\mathcal{O}(10^{-11})$, $\nu_{R}$ will never enter the thermal bath, hence, thermal $\Delta N_{\rm eff}$ is not possible. On the other hand, $\nu_{R}$ produced non-thermally from the decay of either $\chi$ or $N$ gives rise to negligible contribution to $\Delta N_{\rm eff}$. As mentioned before, the Case-B, with $0.01\leq y_{1}^{\prime}\leq 0.1$, $0.03\leq y_{2}^{\prime}\leq 0.1$, $0.03\leq y_{3}^{\prime}\leq 0.1$, we have thermal contribution to $\Delta N_{\rm eff}$. Due to the decoupling of $\nu_{R}$ much above the electroweak scale, we have a fixed contribution $\Delta N_{\rm eff}=0.14$ for Case-B. Finally, in Case-C, Yukawa couplings are of the size $\mathcal{O}(10^{-6})$, and the corresponding interactions do not allow $\nu_{R}$ to enter the thermal bath. Considering the decay of $\chi\rightarrow\nu_{R}N$, we have shown the non-thermal contributions to $\Delta N_{\rm eff}$ in Fig. 10. Clearly, some part of the parameter space is already ruled out by PLANCK 2018 and DESI 2024 limits while most of the remaining parameter space remains within reach of future CMB experiments like SPT-3G, CMB-S4 and CMB-HD. As there are multiple $\mathcal{Z}_{2}$-odd scalars, it is possible to have $\chi_{i}\rightarrow\nu_{R}N$ type decays from one of the heavier scalars irrespective of fermion or scalar DM. While contribution to $\Delta N_{\rm eff}$ from scalar decay can be calculated simply by considering the scalar $\chi_{i}$ to be in equilibrium, contributions from the decay $N\rightarrow\nu_{R}\chi_{j}$ will be more involved as the number density of non-thermal $N$ needs to be found by solving the respective Boltzmann equation Biswas:2022vkq .
@@ VI Other detection prospects
Firstly, we can have a lepton flavour violating (LFV) process at one-loop level, as shown in Fig. 11. The corresponding decay width can be calculated using the prescriptions given in Kuno:1999jp ; Guo:2020qin . With relevant Yukawa couplings of $y_{i}\sim\mathcal{O}(10^{-1}-10^{-2})$, as in Case-A, it gives the branching ratio $\text{Br}(\mu\rightarrow e\gamma)\sim\mathcal{O}(10^{-15})$, which is close to the recent upper bound indicated by MEG-II experiment $\text{Br}(\mu\rightarrow e\gamma)<3.1\times 10^{-13}$ MEGII:2023ltw . For Case-B and Case-C with smaller Yukawa couplings $y_{i}\sim\mathcal{O}(10^{-11})$ and $y_{i}\sim\mathcal{O}(10^{-7})$, we get very small branching ratio $\text{Br}(\mu\rightarrow e\gamma)$ as $\mathcal{O}(10^{-57})$ and $\mathcal{O}(10^{-41})$ respectively, beyond experimental reach in near future.
@@
Due to the presence of several new scalars, admixtures of singlet and doublet, we can have different collider prospects as well. If scalar DM mass $(M_{\chi})$ is lighter than the SM Higgs mass $(M_{h})$ such that $h\rightarrow\chi\chi$ decay is allowed, it can contribute to invisible Higgs decay. Such invisible Higgs decay is currently constrained at Br$(h\rightarrow{\rm invisible})<0.14$ ATLAS:2022yvh . Production of $\mathcal{Z}_{2}$-odd scalars in colliders like the large hadron collider (LHC) can give rise to pure leptonic final states plus missing transverse energy (MET) Gustafsson:2012aj ; Datta:2016nfz , hadronic final states plus MET Poulose:2016lvz or a mixture of both. Such MET can arise from light Dirac neutrinos or scalar/fermion DM in our scenario. It is also possible to have more exotic signatures like tri-lepton plus MET Hashemi:2016wup and mono-jet Belyaev:2016lok ; Belyaev:2018ext . It should be noted that we also have new scalars in the model which are $\mathcal{Z}_{2}$-even. They can lead to additional final states, similar to generic two Higgs doublet models. The new charged scalars can also modify the diphoton decay width of the SM Higgs, currently constrained at $\frac{{\rm BR}(h\rightarrow\gamma\gamma)_{\rm expt}}{{\rm BR}(h\rightarrow\gamma\gamma)_{\rm SM}}=1.12\pm 0.09$ CMS:2021kom .
@@
Another interesting detection prospect arises in the gravitational wave (GW) frontier. Spontaneous breaking of discrete symmetries like $A_{4}\times Z_{2}$, as we have in our model, can lead to formation of topological defects like domain walls (DW). While DW, if stable or allowed to dominate the energy density of the Universe, will be in conflict with cosmological observations, they can be made to disappear by introducing small bias, in the form of explicit symmetry breaking terms. Such annihilating or collapsing DW then generates stochastic GW which can be probed at ongoing or future experiments. GW signatures of minimal Dirac seesaw model was studied in earlier works Barman:2022yos ; Barman:2023fad while similar studies in the context of flavour symmetric models like $A_{4}$ can be found in Gelmini:2020bqg . The bias or explicit symmetry breaking terms introduced for DW disappearance may also lead to interesting indirect detection signatures of DM due to its decay into different standard model particles [^2] See King:2023ayw ; Borah:2024kfn for related discussions.. While the model has several detection prospects, it can also be falsified by future observation of neutrinoless double beta decay.